# Do Determinants of EV Purchase Intent vary across the Spectrum? Evidence from Bayesian Analysis of US Survey Data

### v2

@@ Abstract
journal: Transportation Research Part A
@@ 1 Introduction
Transportation is the largest source of greenhouse gas emissions in the United States (US), representing 28% of the total emissions in 2022, with light-duty vehicles responsible for more than half (Environmental Protection Agency, 2024). Electric vehicles (EVs), given their zero tailpipe emissions, are widely regarded as a key solution for reducing emissions and meeting climate targets (Bhat et al., 2025). To accelerate EV adoption, governments worldwide, including the US, have implemented a range of policies and incentives (Jenn et al., 2018; Mekky and Collins, 2024). Key US measures have included nationwide greenhouse gas standards to reduce emissions from new vehicles; up to \$7,500 federal tax credits for EV buyers; investments in charging infrastructure; and state-level mandates like California’s Zero-Emission Vehicle (ZEV) program, which requires automakers to sell an increasing share of EVs [^1] For more details, see the EPA’s greenhouse gas regulations, the IRS EV tax credit guidelines, National Electric Vehicle Infrastructure (NEVI) Program and California’s ZEV Program. As of late 2025, several of these EV-focused policies have been rolled back under the Trump administration..
@@
Despite these efforts, US sales remain relatively low, accounting for only 10% of global new EV sales in 2023, trailing behind China (60%) and Europe (25%) (International Energy Agency, 2024). The trajectory of US adoption is also increasingly uncertain. In 2024, several automakers, including GM and Ford, scaled back ambitious EV targets, citing weak consumer demand (Colias and Otts, 2024). Successive administrations have taken sharply different approaches to EVs, creating uncertainty around long-term incentives and regulations. The Biden administration promoted EV adoption through subsidies, infrastructure investments, and emissions regulations as part of its broader climate agenda. In contrast, the Trump administration has pursued a deregulatory shift: as of late 2025, the federal EV tax credits are being phased out and the EPA’s endangerment finding is under reconsideration, and California’s Clean Air Act waiver for state-level ZEV mandates has been revoked. [^2] For details on the EV tax credit phase-out, see Section 70502 of the One Big Beautiful Bill Act. For details on the status of California’s waiver for ZEV mandates, see H.J. Res. 88. For details on the reconsideration of the endangerment finding, see the EPA website. Federal charging-infrastructure funding was also suspended but is slated for restoration following litigation (see Department of Transportation newsroom). Other policy changes are subject to ongoing or possible future legal challenges, so further changes remain possible. These changing political priorities reflect tensions between environmental regulations, industry interests, and fiscal considerations in promoting EV adoption. Although EVs reduce emissions—particularly in regions with cleaner electricity grids—subsidies and incentives, along with declining gas tax revenue, pose budgetary challenges (Banta et al., 2024). The Trump administration has also cited concerns that stringent emissions and fuel economy regulations increase compliance costs for automakers and limit consumer choice.
Ultimately, the adoption of EVs depends on consumer interest, which is likely to be shaped by multiple factors, such as cost, convenience, perception of environmental benefits, and awareness of these trade-offs (Rezvani et al., 2015). Although EVs often have higher upfront costs due to expensive batteries, they can offer long-term savings as electric motors require less maintenance (Liu et al., 2021). However, the overall cost of ownership can vary significantly depending on factors such as local gasoline and electricity prices, climate, purchase incentives, home charging access, and annual mileage (Woody et al., 2024). Access to reliable charging infrastructure also remains an important consideration (Chakraborty et al., 2019, 2022). Limited availability in some regions can exacerbate range anxiety and inconvenience, particularly for long-distance travel. Environmental concerns also affect consumer attitudes, but perceptions may vary, given the regional differences in short term benefits (Holland et al., 2016) and the perception about critical minerals mining (Lakshman, 2024).
Recognizing the importance of public interest in shaping the future of this market, we examine how stated purchase intent in the US relates to the following four key factors: exposure to EV information, perceptions of their environmental benefits (compared to gas-powered vehicles), perceptions of government action to mitigate climate change, and confidence in infrastructure development. At the same time, we control for prior ownership, and commonly used demographic variables such as age, gender, education, political affiliation, and geographic location. The data for these four focal factors, along with the dependent and control variables, are drawn from three nationally representative opinion surveys conducted by the Pew Research Center over multiple years.
Our empirical analysis consists of two stages. First, we estimate an ordinal probit model using Markov chain Monte Carlo (MCMC) methods to assess how these factors relate to an ordinal measure of EV purchase intent. We present the covariate effects on the average probabilities of reporting each level of interest, an aspect rarely explored in the EV literature. Secondly, and perhaps more importantly, we extend the analysis using ordinal quantile models at the 20th, 50th, and 80th percentiles of purchase intent to examine how these factors shape the decisions of consumers who are more inclined to adopt EVs compared to those who are less inclined. The results reveal considerable heterogeneity in covariate effects across the distribution of EV purchase intent, which would be missed by ordinal probit models.
@@
Our paper makes three key contributions. First, we examine four unique factors (while controlling for a host of demographic variables) to understand how informational, perceptual, and trust-related factors influence the likelihood and intensity of stated EV purchase intent in the US, and how do these effects vary across different levels of consumer interest. Prior studies of EV purchase intent have considered related factors, but these have either been distinct from the ones we study or investigated only in small samples. The factors we examine have not been analyzed jointly, on large samples, or in the US context. The first factor is confidence in future charging infrastructure development. While earlier work has emphasized the role of existing infrastructure (Carley et al., 2013; Tiwari et al., 2020), purchase intent is also likely to depend on expectations about future network expansion, particularly when the current system remains incomplete. The second factor is perceptions of government climate efforts. Past studies have examined links between purchase intent and political affiliation (e.g., Mamkhezri (2025); Sintov et al. (2020)), but not views on government climate action. Because EVs are widely recognized as a climate policy instrument, individuals who believe the government is already doing enough may be less inclined to purchase—especially if they are less aware of other benefits.
@@
The third factor is exposure to EV-related information. Previous research has examined the effects of information on total cost of ownership (Dumortier et al., 2015), media coverage (Scherrer, 2023), and peer effects (Zhao et al., 2022). We build on this by using a nationally representative survey that directly asks respondents how much they have heard or read about EVs; although subjective, this measure reduces problems of indirect inference. The fourth factor is perceptions of EVs’ environmental benefits. Several studies have considered consumers’ pro-environmental attitudes (Hu et al., 2023; Carley et al., 2013; Krause et al., 2013; Khatua et al., 2023; Scherrer, 2023; Mamkhezri, 2025) and concern about reputation (Buhmann and Criado, 2023). However, research on perceptions of EV’s environmental benefits relative to gasoline powered vehicles is limited. Lashari et al. (2021) and Jia et al. (2025) examine this factor in a small sample, whereas our analysis situates it in a broader US context. [^3] Studies have also examined other factors such as range anxiety, performance concerns, and consumer incentives (See Carley et al. (2013); Tiwari et al. (2020); Zhao et al. (2022); Stekelberg and Vance (2024); Mamkhezri (2025) among others). A separate strand of research examines determinants of actual EV sales, including tax credits and rebates (DeShazo et al., 2017; Jenn et al., 2018; Clinton and Steinberg, 2019; Sheldon et al., 2023), charging infrastructure (Li et al., 2017), peer effects (Chakraborty et al., 2022), and automakers’ responses to environmental policies (Aghion et al., 2016; Gillingham, 2022; Lohawala, 2023). Our findings reveal sizable gaps between the importance of these factors and their prevalence: both infrastructure confidence and information exposure remain low despite their strong association with purchase intent, highlighting opportunities for targeted outreach and investment.
@@
Our second contribution is the use of large, nationally representative datasets, which stands in contrast to the predominantly small-scale, non-representative surveys commonly used in prior research on EV purchase intent. For example, Tiwari et al. (2020), Hu et al. (2023), and Mamkhezri (2025) analyzed 1,800 responses from UK residents, 807 responses from China, and 1,500 responses from the United States, respectively. [^4] Although some recent studies (e.g., Pani et al. (2023) and Ruan and Lv (2022)) use large-scale online sentiment analysis to track the evolution of public perceptions of EVs, they do not examine individual-level predictors of it. Singh et al. (2020) provide a meta-analysis of 211 peer-reviewed research articles 2009–2019, identifying key factors influencing public opinion of EVs. Notably, most such studies are based on single-year, non-representative, or small samples. We overcome these limitations by using three nationally representative US opinion polls, with significantly larger samples: 11,052 in 2021; 7,173 in 2022; and 7,201 in 2023. This allows for greater statistical power and more robust generalizability of our findings.
Our third-and perhaps most significant-contribution lies in applying advanced modeling and estimation techniques to examine the full spectrum of EV purchase intent, addressing the previously unexplored question of whether the determinants of EV adoption vary across different levels of intent. Research has used various methods, such as linear regression to assess the effects of personal beliefs, preferences (Carley et al., 2013; Krause et al., 2013) and positive media coverage (Scherrer, 2023); binary probit and logit to explore motivational factors like reputation (Buhmann and Criado, 2023); mixed logit models to explore attitudes toward various EV policy incentives, employment changes, and electricity costs (Mamkhezri, 2025); structural equation modeling to examine the role of adoption barriers, such as range anxiety and charging infrastructure reliability (Tiwari et al., 2020), peer effects (Zhao et al., 2022), and policy incentives (Hu et al., 2023); and topic modeling and sentiment analysis to track shifts in online sentiment toward EVs over time (Ruan and Lv, 2022). A key limitation of these approaches is that they cannot capture heterogeneity in covariate effects across the full distribution of intent. We address this by applying ordinal quantile analysis to EV purchase intent, an approach not previously used in this context. In contrast to prior work that focuses on average effects, this approach captures heterogeneity in covariate effects both across the full distribution of the intent and across the observed ordinal outcome categories (ranging from “not at all likely” to “very likely” to purchase). This yields more detailed evidence on how different factors shape adoption intent. For example, it shows that confidence in infrastructure development and belief in environmental benefits are not only associated with higher intent among likely adopters, but also with a lower probability of reporting no interest among more skeptical respondents. Note that no classical estimation methods exist for ordinal quantile models, so we use a Bayesian approach.
Finally, our use of a Bayesian framework provides several advantages over traditional (frequentist) methods, which provide point estimates and confidence intervals, but fail to capture uncertainty in model parameters. This can bias covariate effects, particularly in non-linear models like probit or logit, where parameters do not directly reflect changes in outcome probabilities (Jeliazkov and Vossmeyer, 2018). In contrast, Bayesian framework allows us to obtain posterior distribution over the parameters, systematically incorporates uncertainty in parameters (and observations) when computing covariate effects, and enables making direct probability statements about the parameters-rather than relying on the notion of statistical confidence as in the classical framework.
@@ 2 Data
The data utilized offers at least three distinct advantages over existing studies that use survey data to gauge EV purchase intent. First, they are nationally representative and span multiple years. As a result, the estimates we generate are more likely to be robust to cross-sectional heterogeneity. Second, the availability of ordinal responses to likelihood of purchasing EV enables us to model public preference regarding EV purchase in much greater detail as compared to studies that use binary responses, such as (Buhmann and Criado, 2023). Third, the first and second surveys correspond to the Covid-19 period, while the third survey was conducted in the post-pandemic phase. Therefore, if there was an increase in interest for sustainable products like electric vehicles during the pandemic, this should be evident in the data summary, with a higher percentage of respondents expressing a preference for buying electric vehicles during the pandemic phase.
@@ 3.1 The Ordinal Probit Model
Ordinal data models (Johnson and Albert, 2000), also known as ordered choice models, allows fitting a regression model to an ordinal dependent (outcome or response) variable, typically denoted by $y$, as a function of the covariates. These models are grounded in the random utility framework from economics and explained, amongst others, in Jeliazkov and Rahman (2012) and Batham et al. (2023). In a typical setting, the response variable $y$ has several categories and each category is assigned a score (value or number) which is inherently ordered or ranked, without any cardinal interpretation. For example, in our study, responses ($y$) to the question about purchasing an EV are coded as follows: 1 for ‘not at all likely’, 2 for ‘not too likely’, 3 for ‘somewhat likely’, and 4 for ‘very likely’. Here, a score of 2 as compared to 1 implies more inclination to buy an electric vehicle, but we cannot interpret a score of 2 as twice the inclination compared to a score of 1. Similarly, the difference in inclination between the scores 2 and 1 is not the same as that between 4 and 3.
The ordinal regression model, following Albert and Chib (1993), can be written in terms of a continuous latent variable $z_{i}$ as follows:
@@
where $\infty=\gamma_{0}<\gamma_{1}\cdots<\gamma_{J-1}<\gamma_{J}=\infty$ are the cut-points (or thresholds) and $J$ denotes the number of categories or outcomes of $y$. This relationship between the latent $z$ (intent to purchase an EV) and observed $y$ (response to the question about purchasing an EV) is illustrated in Figure 2. As shown, $P(y_{i}=1)$ is the area under $f(z_{i})$ to the left of $\gamma_{1}$, $P(y_{i}=2)$ is the area under $f(z_{i})$ between $\gamma_{1}$ and $\gamma_{2}$, and so forth. To specify the location and scale restrictions, we fix $\gamma_{1}=0$ and $var{(\varepsilon)}=1$, respectively (Jeliazkov et al., 2008; Jeliazkov and Rahman, 2012).
@@
where $\Phi(\cdot)$ is the cumulative distribution function (cdf) of a standard normal distribution and $I(y_{i}=j)$ is an indicator function, which equals 1 if the condition within parenthesis is true and 0 otherwise. But the ordering constraint $\gamma_{1}=0<\gamma_{2}<...<\gamma_{J-1}$ is difficult to satisfy during the MCMC sampling. So, following Albert and Chib (2001), we apply the transformation $\delta_{j}=\ln(\gamma_{j}-\gamma_{j-1})$, for $2\leq j\leq J-1$, and rewrite the likelihood as a function of $\beta$ and $\delta=(\delta_{2},\ldots,\delta_{J-1})^{\prime}$, which allows us to make inferences from $f(y|\beta,\delta)$.
@@
$\begin{split}\pi\left(\beta,\delta,z|y\right)&\propto f\left(y|\beta,\delta,z\right)f(z|\beta)\pi(\beta)\pi(\delta)=\left\{\prod\limits_{i=1}^{n}f\left(y_{i}|\delta,z_{i}\right)\right\}f(z|\beta)\pi(\beta)\pi(\delta),\end{split}$ (4)
where $f(y_{i}|\delta,z_{i})=1\{\gamma_{j-1}<z_{i}\leq\gamma_{j}\}$, with the index $j$ determined by $y_{i}$, $f(z|\beta)=f_{N}(z|X\beta,I_{n})$, and $\pi(\beta)$ and $\pi(\delta)$ are the normal priors. Given the complete data posterior (4), the conditional posteriors can be derived and the model estimated using MCMC sampling as outlined in Algorithm 1.
Post estimation, an important construct is to compute the covariate effects (or marginal effects) on the outcome probabilities, since, unlike linear regression, the coefficients do not give the covariate effects. Let the $l$-th covariate, $x_{i,l}$}, take two distinct values, $a$ and $b$, denoted as $x_{i,l}^{a}$} and $x_{i,l}^{b}$}, respectively. We partition the covariate and parameter vectors as follows: $x_{i}^{a}=(x_{i,l}^{a},x_{i,-l})$, $x_{i}^{b}=(x_{i,l}^{b},x_{i,-l})$, and $\beta=(\beta_{l},\beta_{-l})$, where l-l denotes all covariates (and parameters) except for the $l$-th covariate (and parameter). We evaluate the distribution of the difference $\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})$} for $1\leq j\leq J$, by marginalizing over $x_{i,-l}$} and $(\beta,\delta)$. This enables us to account for sampling and parameter uncertainty, and avoid nontrivial biases in estimating covariate effects (Jeliazkov and Vossmeyer, 2018).
@@
To obtain draws from the distribution $\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})$}, we use the method of composition (Jeliazkov et al., 2008). This involves selecting an individual, extracting the corresponding covariate values, drawing $(\beta,\delta)$ from their posterior distributions, and evaluating the difference $\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta,\delta)-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta,\delta)$}, where $\Pr(y_{i}=j|x_{i}^{q},\beta,\delta)=\Phi(\gamma_{j}-x_{i,l}^{q}\,\beta_{l}-x^{\prime}_{i,-l}\,\beta_{-l})-\Phi(\gamma_{j-1}-x_{i,l}^{q}\,\beta_{l}-x^{\prime}_{i,-l}\,\beta_{-l}),$ for $q=b,a$, and $1\leq j\leq J$. The process is repeated for all individuals and MCMC draws. The average covariate effect (ACE) for outcome $j$ is then computed as the mean of pointwise probability differences:

### v3

@@ Abstract
journal: Transportation Research Part A: Policy and Practice
@@ 1 Introduction
Transportation is the largest source of greenhouse gas emissions in the United States (US), representing 28% of the total emissions in 2022, with light-duty vehicles responsible for more than half [18]. Electric vehicles (EVs), given their zero tailpipe emissions, are widely regarded as a key solution for reducing emissions and meeting climate targets [6]. Over the past decade, governments worldwide, including the US, have implemented a range of policies and incentives aimed at encouraging EV adoption [28, 43]. Key US measures have included nationwide greenhouse gas standards to reduce emissions from new vehicles; up to \$7,500 federal tax credits for EV buyers; investments in charging infrastructure; and state-level mandates like California’s Zero-Emission Vehicle (ZEV) program, which requires automakers to sell an increasing share of EVs [^1] For more details, see the EPA’s greenhouse gas regulations, the IRS EV tax credit guidelines, National Electric Vehicle Infrastructure (NEVI) Program and California’s ZEV Program. As of early 2026, several of these EV-focused policies have been rolled back under the Trump administration..
@@
Despite these efforts, US sales remain relatively low, accounting for only 10% of global new EV sales in 2023, trailing behind China (60%) and Europe (25%) [23]. The trajectory of US adoption is also increasingly uncertain. In 2024, several automakers, including GM and Ford, scaled back ambitious EV targets, citing weak consumer demand [15]. Successive administrations have taken sharply different approaches to EVs, creating uncertainty around long-term incentives and regulations. The Biden administration promoted EV adoption through subsidies, infrastructure investments, and emissions regulations as part of its broader climate agenda. In contrast, the Trump administration has pursued a deregulatory shift: as of early 2026, federal EV tax credits have been eliminated, the EPA’s endangerment finding is under reconsideration, and California’s Clean Air Act waiver for state-level ZEV mandates has been revoked. [^2] For details on the EV tax credit phase-out, see Section 70502 of the One Big Beautiful Bill Act. For details on the status of California’s waiver for ZEV mandates, see H.J. Res. 88. For details on the reconsideration of the endangerment finding, see the EPA website. Federal charging-infrastructure funding was also suspended but is slated for restoration following litigation (see Department of Transportation newsroom). Other policy changes are subject to ongoing or possible future legal challenges, so further changes remain possible. These changing political priorities reflect tensions between environmental regulations, industry interests, and fiscal considerations in promoting EV adoption. Although EVs reduce emissions—particularly in regions with cleaner electricity grids—subsidies and incentives, along with declining gas tax revenue, pose budgetary challenges [4]. The Trump administration has also cited concerns that stringent emissions and fuel economy regulations increase compliance costs for automakers and limit consumer choice.
Amid fluctuating incentives and persistent regulatory uncertainty, EV adoption ultimately depends on consumer interest. Accordingly, the study of public opinion on EV adoption is particularly important, as public opinion is known to influence US public policy even in the presence of competing interests and partisan conflict [9, 50]. While the literature on EV adoption is extensive, important knowledge gaps remain. We organize these gaps, and our contributions to addressing them, along three dimensions: (a) the covariates (independent variables or regressors) considered, (b) the scale and representativeness of the data, and (c) the statistical models used to examine the relationship between EV adoption and these covariates.
Within the first dimension, EV adoption is typically modeled as a function of demographic characteristics–such as age, gender, education, race, political affiliation, and income–together with factors related to cost [39, 56], convenience and charging infrastructure [12, 11], perception of environmental benefits [20, 35], and awareness of these trade-offs [47]. In this context, any determinant that plausibly affects EV adoption but has not been systematically examined represents an opportunity to deepen our understanding of adoption behavior. Building on this literature, we examine EV adoption using standard demographic controls and introduce four additional factors: confidence in building charging stations and infrastructure, perception of government climate efforts, exposure to EV-related information, and perceptions of EVs’ environmental benefits relative to gasoline powered vehicles.
@@
The first factor–confidence in development of future charging infrastructure–is crucial for EV adoption because purchase intent is likely to depend on expectations about future network expansion, particularly when the current system remains incomplete. This stands in contrast to existing studies such as Carley et al. [10] and Tiwari et al. [55], which emphasize the role of existing infrastructure. The second factor concerns perceptions of government climate efforts. While previous research has documented links between EV purchase intent and political affiliation (e.g., Mamkhezri [42], Sintov et al. [53]), it has not paid attention to individuals’ assessments of government climate action itself. Because EVs are widely recognized as a climate policy instrument, individuals who believe the government is already doing enough may be less inclined to purchase an EV—particularly if they are less aware of EVs’ ancillary benefits.
@@
The third factor is exposure to EV-related information. Previous research has examined the effects of information on total cost of ownership [17], media coverage [49], and peer effects [58]. We extend these works by using a nationally representative survey that directly asks respondents how much they have heard or read about EVs; although subjective, this measure reduces problems of indirect inference. The fourth factor is perceptions of EVs’ environmental benefits. Several studies have considered consumers’ pro-environmental attitudes [22, 10, 33, 31, 49, 42] and reputational concerns [8]. However, research on perceptions of EV’s environmental benefits relative to gasoline powered vehicles is limited. Lashari et al. [36] and Jia et al. [29] examine this factor in a small sample, whereas our analysis situates it in a broader US context. [^3] Studies have also examined other factors such as range anxiety, performance concerns, and consumer incentives (See Carley et al. [10], Tiwari et al. [55], Zhao et al. [58], Stekelberg and Vance [54], Mamkhezri [42] among others). A separate strand of research examines determinants of actual EV sales, including tax credits and rebates [16, 28, 14, 51], charging infrastructure [37], peer effects [11], and automakers’ responses to environmental policies [1, 19, 40].
@@
Turning to the second dimension–the scale and representativeness of data–we observe that much of the existing research on EV purchase intent relies on single-year surveys, non-representative samples, or relatively small datasets. For example, Tiwari et al. [55], Hu et al. [22], and Mamkhezri [42] analyzed 1,800 responses from UK residents, 807 responses from China, and 1,500 responses from the US, respectively. [^4] Although some recent studies (e.g., Pani et al. [44] and Ruan and Lv [48]) use large-scale online sentiment analysis to track the evolution of public perceptions of EVs, they do not examine individual-level predictors of it. Singh et al. [52] provide a meta-analysis of 211 peer-reviewed research articles between 2009–2019, identifying key factors influencing public opinion of EVs. We overcome the limitations associated with small data by using three nationally representative US opinion polls, with substantially larger samples: 11,052 in 2021; 7,173 in 2022; and 7,201 in 2023. This allows for greater statistical power and more robust generalizability of our findings
Finally, along the third dimension–use of statistical models to analyze EV adoption–we note that most studies have relied on linear regression to assess the effects of personal beliefs, preferences [10, 33] and positive media coverage [49]. Some studies have employed binary probit and logit models to examine motivational factors such as reputation [8]; while others have used mixed logit models to analyze attitudes toward various EV policy incentives, employment changes, and electricity costs [42]. However, in much of this literature, covariate effects surprisingly remains unreported, limiting the interpretability of the estimated relationships. Beyond these approaches, structural equation models have also been used to examine adoption barriers [55], peer effects [58], and policy incentives [22]; while topic modeling and sentiment analysis have been applied to track temporal shifts in online sentiment toward EVs [48]. Nonetheless, all the above mentioned modeling frameworks cannot capture the heterogeneity in covariate effects across the full distribution of intent, which is necessary for a careful and deeper understanding of EV adoption.
We employ a Bayesian approach and contribute to the modeling of EV adoption literature in two key ways. First, we estimate an ordinal probit model using Markov chain Monte Carlo (MCMC) methods to assess how our four factors and demographic characteristics relate to an ordinal measure of EV purchase intent. We report covariate effects on the average probabilities of each response category, an aspect that has received limited attention in prior studies. Second, and more importantly, we apply ordinal quantile analysis to examine the full spectrum of EV purchase intent, addressing the previously unexplored question of whether the determinants of EV adoption vary across different levels of intent–a question that cannot be answered using conditional mean models (e.g., ordinal probit). We note that this is the first application of ordinal quantile modeling in the EV adoption literature, enabling us to capture heterogeneity in covariate effects both across the full distribution of the intent and across the observed ordinal outcome categories (ranging from “not at all likely” to “very likely” to purchase).
The results from ordinal probit models indicate that greater exposure to EV-related information, stronger perceptions of EVs’ environmental benefits, and higher confidence in future charging infrastructure development are positively associated with EV purchase intent, whereas perceptions of stronger government climate action are negatively associated with intent across samples. These relationships persist after controlling for standard demographic characteristics: younger, urban, more educated, and higher-income individuals are more likely to express interest in EVs, while women, US-born, and Black respondents report lower intent. The results from ordinal quantile analysis uncover substantial heterogeneity in covariate effects across the distribution of EV purchase intent. This approach provides richer insight into how different factors shape adoption intent. For example, it shows that confidence in infrastructure development and belief in environmental benefits are not only associated with higher intent among likely adopters, but also with a lower probability of reporting no interest among more skeptical respondents. Finally, we identify notable gaps between the importance of these factors and their prevalence: both infrastructure confidence and information exposure remain low despite their strong association with purchase intent, highlighting opportunities for targeted outreach and investment.
We conclude this section by motivating our choice of a Bayesian approach for estimating the ordinal models. In non-linear models such as the ordinal probit, regression coefficients do not directly correspond to covariate effects–defined as changes in outcome probabilities–and these effects must therefore be computed separately for each outcome category. Proper estimation of covariate effects requires accounting for uncertainty in both observed variables and model parameters. But the latter is not captured in the Classical framework resulting in bias in covariate effects [27]. In contrast, the Bayesian framework provides a full posterior distribution over the parameters and allows uncertainty to be systematically incorporated when computing covariate effects. Moreover, our adoption of Bayesian methods is driven by methodological necessity: to the best of our knowledge, frequentist estimation procedures for ordinal quantile models are not currently available, making Bayesian estimation essential in this context.
@@ 2 Data
The data utilized offers at least three distinct advantages over existing studies that use survey data to gauge EV purchase intent. First, they are nationally representative and span multiple years. As a result, the estimates we generate are more likely to be robust to cross-sectional heterogeneity. Second, the availability of ordinal responses to likelihood of purchasing EV enables us to model public preference regarding EV purchase in much greater detail as compared to studies that use binary responses, such as [8]. Third, the first and second surveys correspond to the Covid-19 period, while the third survey was conducted in the post-pandemic phase. Therefore, if there was an increase in interest for sustainable products like electric vehicles during the pandemic, this should be evident in the data summary, with a higher percentage of respondents expressing a preference for buying electric vehicles during the pandemic phase.
@@ 3.1 The Ordinal Probit Model
Ordinal data models [30], also known as ordered choice models, allows fitting a regression model to an ordinal dependent (outcome or response) variable, typically denoted by $y$, as a function of the covariates. These models are grounded in the random utility framework from economics and explained, amongst others, in Jeliazkov and Rahman [26] and Batham et al. [5]. In a typical setting, the response variable $y$ has several categories and each category is assigned a score (value or number) which is inherently ordered or ranked, without any cardinal interpretation. For example, in our study, responses ($y$) to the question about purchasing an EV are coded as follows: 1 for ‘not at all likely’, 2 for ‘not too likely’, 3 for ‘somewhat likely’, and 4 for ‘very likely’. Here, a score of 2 as compared to 1 implies more inclination to buy an electric vehicle, but we cannot interpret a score of 2 as twice the inclination compared to a score of 1. Similarly, the difference in inclination between the scores 2 and 1 is not the same as that between 4 and 3.
The ordinal regression model, following Albert and Chib [2], can be written in terms of a continuous latent variable $z_{i}$ as follows:
@@
where $\infty=\gamma_{0}<\gamma_{1}\cdots<\gamma_{J-1}<\gamma_{J}=\infty$ are the cut-points (or thresholds) and $J$ denotes the number of categories or outcomes of $y$. This relationship between the latent $z$ (intent to purchase an EV) and observed $y$ (response to the question about purchasing an EV) is illustrated in Figure 2. As shown, $P(y_{i}=1)$ is the area under $f(z_{i})$ to the left of $\gamma_{1}$, $P(y_{i}=2)$ is the area under $f(z_{i})$ between $\gamma_{1}$ and $\gamma_{2}$, and so forth. To specify the location and scale restrictions, we fix $\gamma_{1}=0$ and $var{(\varepsilon)}=1$, respectively [25, 26].
@@
where $\Phi(\cdot)$ is the cumulative distribution function (cdf) of a standard normal distribution and $I(y_{i}=j)$ is an indicator function, which equals 1 if the condition within parenthesis is true and 0 otherwise. But the ordering constraint $\gamma_{1}=0<\gamma_{2}<...<\gamma_{J-1}$ is difficult to satisfy during the MCMC sampling. So, following Albert and Chib [3], we apply the transformation $\delta_{j}=\ln(\gamma_{j}-\gamma_{j-1})$, for $2\leq j\leq J-1$, and rewrite the likelihood as a function of $\beta$ and $\delta=(\delta_{2},\ldots,\delta_{J-1})^{\prime}$, which allows us to make inferences from $f(y|\beta,\delta)$.
@@
$\begin{split}\pi\left(\beta,\delta,z|y\right)&\propto f\left(y|\beta,\delta,z\right)f(z|\beta)\pi(\beta)\pi(\delta)=\left\{\prod\limits_{i=1}^{n}f\left(y_{i}|\delta,z_{i}\right)\right\}f_{N}(z|X\beta,I_{n})\pi(\beta)\pi(\delta),\end{split}$ (4)
where $f(y_{i}|\delta,z_{i})=1\{\gamma_{j-1}<z_{i}\leq\gamma_{j}\}$, with the index $j$ determined by $y_{i}$, and $\pi(\beta)$ and $\pi(\delta)$ are the normal priors. Given the complete data posterior (4), the conditional posteriors can be derived and the model estimated using MCMC sampling as outlined in Algorithm 1.
Post estimation, an important construct is to compute the covariate effects (or marginal effects) on the outcome probabilities, since, unlike linear regression, the coefficients do not give the covariate effects. Let the $l$-th covariate, $x_{i,l}$}, take two distinct values, $a$ and $b$, denoted as $x_{i,l}^{a}$} and $x_{i,l}^{b}$}, respectively. We partition the covariate and parameter vectors as follows: $x_{i}^{a}=(x_{i,l}^{a},x_{i,-l})$, $x_{i}^{b}=(x_{i,l}^{b},x_{i,-l})$, and $\beta=(\beta_{l},\beta_{-l})$, where l-l denotes all covariates (and parameters) except for the $l$-th covariate (and parameter). We evaluate the distribution of the difference $\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})$} for $1\leq j\leq J$, by marginalizing over $x_{i,-l}$} and $(\beta,\delta)$. This enables us to account for sampling and parameter uncertainty, and avoid nontrivial biases in estimating covariate effects [27].
@@
To obtain draws from the distribution $\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})$}, we use the method of composition [25]. This involves selecting an individual, extracting the corresponding covariate values, drawing $(\beta,\delta)$ from their posterior distributions, and evaluating the difference $\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta,\delta)-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta,\delta)$}, where $\Pr(y_{i}=j|x_{i}^{q},\beta,\delta)=\Phi(\gamma_{j}-x_{i,l}^{q}\,\beta_{l}-x^{\prime}_{i,-l}\,\beta_{-l})-\Phi(\gamma_{j-1}-x_{i,l}^{q}\,\beta_{l}-x^{\prime}_{i,-l}\,\beta_{-l}),$ for $q=b,a$, and $1\leq j\leq J$. The process is repeated for all individuals and MCMC draws. The average covariate effect (ACE) for outcome $j$ is then computed as the mean of pointwise probability differences: